At last week’s hockey game involving Waterloo Blueberries, the price of a
platinum ticket was $25, the price of a gold ticket was $10, the price of a silver ticket was $5, and the price of a bronze ticket was $1.
(a) Jim buys 5 platinum tickets, 2 gold tickets and 3 silver tickets. Determine the average cost of the tickets that Jim buys.
(b) Mike buys 8 tickets whose average cost is $12. He then buys five more platinum tickets. What is the new average cost of the tickets that he has bought?
(c) Ophelia buys 10 tickets with an average cost of $14. Suppose that she buys n more platinum tickets. The new average cost of the tickets that she has bought is $20. What is the value of n?

Answer by mananth(16946)

At last week’s hockey game involving Waterloo Blueberries, the price of a
platinum ticket was $25, the price of a gold ticket was $10, the price of a silver ticket was $5, and the price of a bronze ticket was $1.
..
(a) Jim buys 5 platinum tickets,5*25= 125
2 gold tickets 2*10 =20
and 3 silver tickets. 3*5=15
150+20+15= 160
Number of tickets = 10
Average price = 160/10 = 16 $
..
(b) Mike buys 8 tickets whose average cost is $12. He then buys five more platinum tickets. What is the new average cost of the tickets that he has bought?
8 tickets average cost = $12
Total cost =$96
5 platinum = 5*25=125
Total = 125+96=$ 221
Total tickets = 13
Average cost = 221/13= $17
..
(c) Ophelia buys 10 tickets with an average cost of $14. Suppose that she buys n more platinum tickets. The new average cost of the tickets that she has bought is $20. What is the value of n?
10 tickets average = $14
Total cost = $140
n more platinum tickets
cost per ticket = $25
cost of n platinum tickets = $25n
Total number of tickets = 10+n
Average cost of tickets =140+25n / n+10 = 20
140+25n = 20(n+10)
140+25n = 20n +200
25n-20n=200-140
5n = 60
n= 12